OpenStudy (anonymous):

The sum of a number and its reciprocal is 5 1/5.

OpenStudy (unklerhaukus):

take a guess

OpenStudy (anonymous):

do I have to convert the mixed fraction to a improper fraction?

OpenStudy (unklerhaukus):

if you can't see the answer staring at you

OpenStudy (lgbasallote):

sadly, i cannot see the answer staring at me either

OpenStudy (anonymous):

\[x^2+1-5\frac{1 }{ 5}=0?\]

OpenStudy (zzr0ck3r):

5 and 1/2 what does and mean? and what is the recipicle of 5?

OpenStudy (anonymous):

1/5

OpenStudy (zzr0ck3r):

so what does and mean?

OpenStudy (anonymous):

but I don't get what the question is asking.. it says what is the reciprocal of 5 and 1/5... but how is it 5 or 1/5?

OpenStudy (zzr0ck3r):

its asking what two numbers can you add together and get 5 and 1/2 or what numbers can you add together and get 5+1/2 note: 5 is the recipicle of 1/5 and vice versa

OpenStudy (lgbasallote):

5 and 1/5*

OpenStudy (zzr0ck3r):

:)

OpenStudy (zzr0ck3r):

5+1/5 = 5 and 1/5 5^(-1) = 1/5

OpenStudy (zzr0ck3r):

hence "the answer is stairing you in the face"

OpenStudy (anonymous):

OH.... so its 5+1/5 not 5 1/5... i was thinking something differently.. thanks

OpenStudy (anonymous):

s othe answer is 5 or 1/5 hahaha

OpenStudy (anonymous):

So there is no real calculation needed for this question?

OpenStudy (unklerhaukus):

\[x+\frac 1x=5~\frac 15\]\[\frac {x^2+1}x=\frac {26}5\]\[x^2+1=\left(\frac {26}5\right)x\]\[x^2-\left(\frac{26}5\right)x+1=0\] \[x_{1,2}=\frac{-(-\frac{26}5)\pm\sqrt{(-\frac{26}5)^2-4}}{2}\]\[\qquad=\frac{\frac{26}5\pm\sqrt{\frac{676}{25}-4}}{2}\]\[\quad=\quad...\]

OpenStudy (anonymous):

x+ 5/26